OpenStudy (anonymous):

Compare each function with the parent function. Without graphing, what are the vertex, axis of symmetry, and transformations of the parent function? y = |6x - 7| - 4

OpenStudy (jdoe0001):

y = A( Bx + C ) + D C = horizontal shift, C > 0, to the left, C < 0, to the right D = vertical shift, D > 0, up, D < 0, down

OpenStudy (anonymous):

So y = A(Bx + C) + D would be the parent function?

OpenStudy (jdoe0001):

the parent function in this case is just y = |x| the other "additions" or numbers are the so-called "transformations" or shifts

OpenStudy (jdoe0001):

B = shrinks or expands the graphic, B > 0, graph opens upwards, B < 0, the graph opens downward

OpenStudy (anonymous):

Since it's absolute value, the graphic would open upwards? |6x|

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

I'm still confused, though. How do I find the vertex?

OpenStudy (jdoe0001):